the ratio of length to width (measured in inches) in a rectangle is 4:7. Find the length of the rectangle if the area is known to be 700 square inches

Answer by stanbon(75887)

the ratio of length to width (measured in inches) in a rectangle is 4:7. Find the length of the rectangle if the area is known to be 700 square inches
-------
4:7 is the same as 4x:7x where 4x is length and 7x is width
-------
Equation:
4x*7x = 700
28x^2 = 700
x^2 = 25
x = 5
-----
length = 4x = 20 inches
width = 7x = 35 inches
--------------
Cheers,
Stan H.
--------------